# Leveraging differentiable programming in the inverse problem of neutron stars

### v1

@@ III.3 Constraints on nuclear empirical parameters
Currently available NS observations do not yet tightly constrain the EOS, as demonstrated by the previous section. In this section, we instead consider a hypothetical scenario where observations of NSs are pushed to extreme limits and uncertainties on the measured radii and tidal deformabilities can be ignored. We aim to obtain a qualitative estimate of how highly accurate knowledge of the properties of NSs translates into constraints on the NEPs, since such constraints can then be compared to those obtained from nuclear experiments. For simplicity, we restrict our parametrization to the metamodel part and no longer use the CSE at high densities. We enforce all these metamodel EOSs to be causal and therefore truncate the EOS once the speed-of-sound reaches $1$. Otherwise, the EOSs are truncated at the maximal central density allowed in an NS.
@@
We generate a mock “true” EOS with a TOV mass of roughly $2.0M_{\odot}$ and aim to recover its EOS parameters from its $M$-$\Lambda$ curve using the gradient-based optimization scheme discussed in Sec. II.6, leaving a full Bayesian inference-based study for future work. We perform multiple such recoveries where we change the number of NEPs that are varied freely in the recovery, starting with only the first order parameter $L_{\rm{sym}}$. In subsequent recovery runs, we also allow the optimizer to explore the higher-order parameters, each time adding the parameters of the next order of the Taylor expansions of Eqs. (2)-(3). For simplicity, we do not vary the $E_{\rm{sym}}$ parameter, since it has a small impact on the EOS at high densities, and we wish to focus on the parameters entering at higher orders of the Taylor expansion. For each recovery, we start the optimizer from $100$ random EOS samples, sampled from the priors shown in Table 1, and perform at most $1000$ iterations with a fixed learning rate of $1$. [^8] The runs are performed on an AMD EPYC 7543 VM CPU and take at most a few minutes per recovery. For the final result, we keep only those EOS samples that within the $[1\ M_{\odot},2\ M_{\odot}]$ mass range deviate less than $100$ meters in radius from the “true” $M$-$R$ curve and less than $10$ in tidal deformability from the “true” $M$-$\Lambda$ curve. Figure 4 shows the resulting recovery on the slope of the symmetry energy $L_{\rm{sym}}$. The range of the recovered $L_{\rm{sym}}$ values significantly depends on the number of NEPs that are allowed to vary. When $L_{\rm{sym}}$ is the only degree of freedom, we recover the value of the “true” EOS, with a small width of around $0.5\ \rm{MeV}$, as expected. However, when more NEPs are varied, this range increases. Nevertheless, the middle and right panels of Fig. 4 demonstrate (for the case where all NEPs are free parameters) that the corresponding EOSs and NS configurations agree well with the true EOS. Therefore, from our gradient-based optimization approach, we conclude that, while accurate observations of NSs allow us to tightly constrain the EOS itself, the corresponding constraints on the NEPs suffer from a degeneracy in our metamodel parametrization, which was already hinted at by Refs. Krastev and Li (2019); Xie and Li (2019, 2020); Mondal and Gulminelli (2022); Iacovelli et al. (2023). Therefore, fitting relations between $L_{\rm{sym}}$ and $R_{1.4}$ previously proposed in the literature, e.g., Ref. Raithel et al. (2021), will need to be assessed carefully in light of the assumptions regarding the parametrization used to obtain them. Finally, we note that this degeneracy is expected to be more pronounced if we include the CSE parametrization at higher densities.
@@
Figure 4: Recovery of an EOS in the metamodel parametrization from the $M$-$\Lambda$ curve using the gradient-based optimization algorithm as a function of the number of NEPs being varied. The left panel shows the recovered range of $L_{\rm{sym}}$ when increasingly more NEPs at higher orders in the Taylor expansion are varied freely. For the recovery in which all NEPs were varied freely, we show the recovered EOS samples in the middle and right panels. The middle panel shows the pressure as a function of density, while the right panel shows the mass-radius profile of NSs, color-coded by their $L_{\rm{sym}}$ value. The grey bands show mass ranges not considered by the loss function in Eq. (13). All EOSs considered show a deviation with respect to the target of $<100$ meters in radius and $<10$ in dimensionless tidal deformabilities across the mass range considered in the loss function of Eq. (13).
@@ IV.1 Equation of state degeneracies
Our results from Sec. III.3 show that the inverse problem of NSs is sensitive to a degeneracy in our metamodel parametrization. While this was shown using our gradient-based optimization scheme, future work will assess if these degeneracies remain when using traditional stochastic sampling algorithms. A similar degeneracy was already observed in previous works that used stochastic samplers, e.g., Refs. Krastev and Li (2019); Xie and Li (2019, 2020); Mondal and Gulminelli (2022); Iacovelli et al. (2023). However, our work extends these studies by simultaneously considering (i) higher-order NEPs, (ii) a wide range of NS masses, (iii) high accuracy in the recovered EOSs, highlighting the problem’s relevance for future experiments, and (iv) the dependence as a function of the parametrization. While our analysis in Sec. III.3 served as a theoretical exploration of this degeneracy, we expect that a full Bayesian analysis employing future constraints from nuclear theory and experiments on the NEPs, especially for the higher-order parameters of the MM, alleviates this issue.
We expect the level of degeneracy to be different for different EOS models, and we will explore such model dependencies in future work. Furthermore, we will also investigate the potential introduction of similar systematic effects originating from certain aspects neglected in this study, such as the crust treatment (see, e.g., Refs. Fortin et al. (2016); Gamba et al. (2020); Ferreira and Providência (2020); Rather et al. (2021); Suleiman et al. (2021); Canullan-Pascual et al. (2025)), and consider unified EOS models Davis et al. (2024) as alternative.
@@ IV.2 Next-generation detectors
Future detectors will generate an unprecedented volume of NS observations. Each event provides valuable constraints on the EOS, yet, when analyzed with existing methods, at substantial economic and environmental costs Hu and Veitch (2024). Our work addresses this pressing data analysis challenge by employing differentiable programming and GPU accelerators. Using jax, analyzing a GW signal from a BNS merger to obtain masses and tidal deformabilities can be completed within minutes Wouters et al. (2024), after which we obtain posteriors on the EOS in less than $1$ hour on an NVIDIA H100 GPU. Crucially, we do not require pre-trained machine-learning emulators of the TOV equations, thereby ensuring accuracy in the results and further reducing computational costs. Therefore, our methods are a key advancement towards handling the data analysis problems offered by future detectors.
Furthermore, while our constraint on the $n_{\rm{break}}$ parameter from Sec. III.2 serves as a proof of principle, we expect that additional observations of NSs will tighten this constraint and allow us to test the consistency between nuclear physics and astrophysics. Future research using simulated data will have to investigate to what accuracy the breakdown density of the metamodel can in fact be constrained from NS data. Moreover, the proof of concept presented here can be extended to directly test predictions from $\chi$EFT, providing an analysis complementary to Ref. Essick et al. (2020a) with an alternative EOS parametrization to assess possible systematic effects in such an analysis.
@@ IV.3 Extension to other analyses
In this work, we only focused on inferring the EOS from NS observations. However, other inferences require to jointly model the EOS together with additional effects or parameters. Naturally, these analyses can therefore benefit from the acceleration offered by the methods presented in this work.
For instance, we can infer the EOS simultaneously with cosmological parameters or population models (see, e.g., Refs. Ghosh et al. (2024); Magnall et al. (2024); Golomb et al. (2025)). Moreover, our method can be extended to jointly infer the parameters of modified gravity, either for a specific theory (see Ref. Olmo et al. (2020) for an overview) or for a parametrized post-TOV formalism Glampedakis et al. (2015, 2016); da Silva et al. (2024). Our gradient-based scheme can be used to explore this extended parameter space and uncover degeneracies between the EOS and deviations from general relativity in observations of NSs Brown et al. (2024). Similar degeneracies can arise from other effects, such as phase transitions Raithel and Most (2023a, b); Counsell et al. (2025) or dark matter Koehn et al. (2024b), which can also be constrained with either Bayesian inference or the gradient-based method.
Going beyond extreme matter studies, other analyses in related fields that have to solve ordinary differential equations on the fly during inference can benefit from the methods discussed in this work, such as the evaluation of effective one-body gravitational waveforms Damour and Nagar (2011), solving the lens equation in strong lensing analyses (see, e.g., Refs. Grespan and Biesiada (2023); Poon et al. (2024)), and nuclear reaction networks Reichert et al. (2023).
@@ V Conclusion
Additionally, our framework introduces a novel optimization scheme that recovers the EOS of a given mass-radius or mass-tidal deformability relation by performing gradient descent on a loss function formulated in terms of the EOS parameters, differentiating through the TOV equations. Using this tool, we have discussed a degeneracy present in our metamodel parametrization, which might affect inferences of nuclear empirical parameters from detailed knowledge of NSs using certain EOS parametrizations.
@@ Appendix A Method validations
In Table 3, we extend the results of Table 2 and show $M_{\rm{TOV}}$, the pressure at $3n_{\rm{sat}}$, $p(3n_{\rm{sat}})$, and the maximal central density $n_{\rm{TOV}}$ at the core of an NS, between Ref. Koehn et al. (2024a) and this work for the EOS constraints discussed in Sec. II.5.
Table 3: Comparison of inferred $M_{\rm{TOV}}$, the pressure at $3n_{\rm{sat}}$, $p(3n_{\rm{sat}})$, and the maximal central density $n_{\rm{TOV}}$ at the core of an NS, between Ref. Koehn et al. (2024a) and this work for various EOS constraints.
Constraint $M_{\rm{TOV}}$ [$M_{\odot}$] $p(3n_{\rm{sat}})$ [MeV fm-3] $n_{\rm{TOV}}$ [$n_{\rm{sat}}$]
Ref. Koehn et al. (2024a) This work Ref. Koehn et al. (2024a) This work Ref. Koehn et al. (2024a) This work
@@
Radio timing $2.35_{-0.29}^{+0.73}$ $2.20_{-0.26}^{+0.39}$ $111_{-49}^{+140}$ $\phantom{0}97_{-49\phantom{0}}^{+65\phantom{0}}$ $5.51_{-1.66}^{+1.89}$ $5.68_{-1.74}^{+1.63}$
@@
All $2.25_{-0.22}^{+0.42}$ $2.23_{-0.23}^{+0.41}$ $\phantom{0}90_{-31\phantom{0}}^{+71\phantom{0}}$ $\phantom{0}92_{-42\phantom{0}}^{+70\phantom{0}}$ $5.92_{-1.38}^{+1.35}$ $5.92_{-1.49}^{+1.38}$
@@
[…]

### v2

@@ III.3 Constraints on nuclear empirical parameters
Currently available NS observations do not yet tightly constrain the EOS, as demonstrated by the previous section. In this section, we instead consider a hypothetical scenario where observations of NSs are pushed to extreme limits and uncertainties on the measured radii and tidal deformabilities can be ignored. We aim to obtain a qualitative estimate of how highly accurate knowledge of the properties of NSs translates into constraints on the NEPs, since such constraints can then be compared to those obtained from nuclear experiments. For simplicity, we restrict our parametrization to the metamodel part and no longer use the CSE at high densities. We enforce all these metamodel EOSs to be thermodynamically stable as well as to respect causality. Otherwise, the EOSs are truncated at the maximal central density allowed in an NS.
@@
We generate a mock “true” EOS [^8] Here, we did not consider agreement between this mock EOS and the constraints from perturbative QCD at high densities, as this merely serves as a proof of concept of the methodology. with a TOV mass of roughly $2.0M_{\odot}$ and aim to recover its EOS parameters from its $M$-$\Lambda$ curve using the gradient-based optimization scheme discussed in Sec. II.6. In particular, the inversion is performed with perfect knowledge of the $M$-$\Lambda$ curve, as discussed in detail in Sec. II.6. We leave a full study which considers Bayesian inference using the constraints from Tab. 2for future work. We perform multiple such recoveries where we change the number of NEPs that are varied freely in the recovery, starting with only the first order parameter $L_{\rm{sym}}$. In subsequent recovery runs, we also allow the optimizer to explore the higher-order parameters, each time adding the parameters of the next order of the Taylor expansions of Eqs. (2)-(3). For simplicity, we do not vary the $E_{\rm{sym}}$ parameter, since it has a small impact on the EOS at high densities, and we wish to focus on the parameters entering at higher orders of the Taylor expansion. For each recovery, we start the optimizer from $100$ random EOS samples, sampled from the priors shown in Table 1, and perform at most $1000$ iterations with a fixed learning rate of 11.999The runs are performed on an AMD EPYC 7543 VM CPU and take at most a few minutes per recovery. For the final result, we keep only those EOS samples that within the $1\ M_{\odot},2\ M_{\odot}]$ mass range deviate less than $100$ meters in radius from the “true” $M$-$R$ curve and less than $10$ in tidal deformability from the “true” $M$-$\Lambda$ curve.
Figure 4 shows the slope of the symmetry energy $L_{\rm{sym}}$ for these recovered EOSs. The range of the recovered $L_{\rm{sym}}$ values significantly depends on the number of NEPs that are allowed to vary. When $L_{\rm{sym}}$ is the only degree of freedom, we recover the value of the “true” EOS, with a small width of around $0.5\ \rm{MeV}$, as expected. However, when more NEPs are varied, this range increases. Nevertheless, the middle and right panels of Fig. 4 demonstrate (for the case where all NEPs are free parameters) that the corresponding EOSs and NS configurations agree well with the true EOS. Therefore, from our gradient-based optimization approach, we conclude that, while accurate observations of NSs allow us to tightly constrain the EOS itself, the corresponding constraints on the NEPs suffer from a degeneracy in our metamodel parametrization, which was already hinted at by Refs. Krastev:2018nwr; Xie:2019sqb; Xie:2020tdo; Mondal:2021vzt; Iacovelli:2023nbv. Therefore, fitting relations between $L_{\rm{sym}}$ and $R_{1.4}$ previously proposed in the literature, e.g., Ref. Raithel:2020vvg, will need to be assessed carefully in light of the assumptions regarding the parametrization used to obtain them. Finally, we note that this degeneracy is expected to be more pronounced if we include the CSE parametrization at higher densities, as this would further diminish the impact of the individual NEPs at the higher-density regime of the EOS to which NSs are most sensitive.
@@
Figure 4: Recovery of an EOS in the metamodel parametrization from the $M$-$\Lambda$ curve using the gradient-based optimization algorithm as a function of the number of NEPs being varied. The left panel shows the recovered range of $L_{\rm{sym}}$ when increasingly more NEPs at higher orders in the Taylor expansion are varied freely. For the recovery in which all NEPs were varied freely, we show the recovered EOS samples in the middle and right panels. The middle panel shows the pressure as a function of density, while the right panel shows the mass-radius profile of NSs, color-coded by their $L_{\rm{sym}}$ value. The grey bands show mass ranges not considered by the loss function in Eq. (13). The EOSs shown in this figure deviate from the target by less than $100$ meters in radius and less than $10$ in dimensionless tidal deformability across the mass range considered in the loss function of Eq. (13).
@@ IV.1 Equation of state degeneracies
Our results from Sec. III.3 show that the inverse problem of NSs is sensitive to degeneracy in our metamodel parametrization. While this was shown using our gradient-based optimization scheme, future work will assess if these degeneracies remain when using traditional stochastic sampling algorithms. A similar degeneracy was already observed in previous works that used stochastic samplers, e.g., Refs. Krastev:2018nwr; Xie:2019sqb; Xie:2020tdo; Mondal:2021vzt; Iacovelli:2023nbv. However, our work extends these studies by simultaneously considering (i) higher-order NEPs, (ii) a wide range of NS masses, (iii) high accuracy in the recovered EOSs, highlighting the problem’s relevance for future detectors, and (iv) the dependence as a function of the parametrization. This degeneracy is expected since the MM expansion from Eq. (2) - (3) is performed at $n=n_{\rm{sat}}$, whereas NSs observations probe the EOS at densities of a few $n_{\rm{sat}}$, hinting that other parameterizations which avoid a Taylor expansion at a specific density could potentially alleviate this issue. While our analysis in Sec. III.3 served as a theoretical exploration of this degeneracy, we expect that a full Bayesian analysis employing future constraints from nuclear theory and experiments on the NEPs, especially for the higher-order parameters of the MM, alleviates this issue.
We expect the level of degeneracy to be different for different EOS models, and we will explore such model dependencies in future work. Furthermore, we will also investigate the potential introduction of similar systematic effects originating from certain aspects neglected in this study, such as the crust treatment (see, e.g., Refs. Fortin:2016hny; Gamba:2019kwu; Ferreira:2020wsf; Rather:2020gja; Suleiman:2021hre; Canullan-Pascual:2025smm), and consider unified EOS models Davis:2024nda as alternative.
@@ IV.2 Next-generation detectors
Future detectors will generate an unprecedented volume of NS observations. Each event provides valuable constraints on the EOS, yet, when analyzed with existing methods, at substantial economic and environmental costs Hu:2024mvn. Our work addresses this pressing data analysis challenge by employing differentiable programming and GPU accelerators. Using jax, analyzing a GW signal from a BNS merger to obtain masses and tidal deformabilities can be completed within minutes Wouters:2024oxj, after which we obtain posteriors on the EOS in less than $1$ hour on an NVIDIA H100 GPU. Crucially, we do not require pre-trained machine-learning emulators of the TOV equations, thereby ensuring accuracy in the results and further reducing computational costs. Therefore, our methods are a key advancement towards handling the data analysis problems offered by future detectors.
[…]
@@ IV.3 Extension to other analyses
In this work, we only focused on inferring the EOS from NS observations. However, other inferences require jointly modeling the EOS together with additional effects or parameters. Naturally, these analyses can therefore benefit from the acceleration offered by the methods presented in this work.
For instance, we can infer the EOS simultaneously with cosmological parameters or population models (see, e.g., Refs. Ghosh:2024cwc; Magnall:2024ffd; Golomb:2024lds). Moreover, our method can be extended to jointly infer the parameters of modified gravity, either for a specific theory (see Ref. Olmo:2019flu for an overview) or for a parametrized post-TOV formalism Glampedakis:2015sua; Glampedakis:2016pes; daSilva:2024qex. Our gradient-based scheme can be used to explore this extended parameter space and uncover degeneracies between the EOS and deviations from general relativity in observations of NSs Brown:2024joy. Similar degeneracies can arise from other effects, such as phase transitions Raithel:2022aee; Raithel:2022efm; Counsell:2025hcv or dark matter Koehn:2024gal, which can also be constrained with either Bayesian inference or the gradient-based method.
Going beyond extreme matter studies, other analyses in related fields that have to solve ordinary differential equations on the fly during inference can benefit from the methods discussed in this work, such as the evaluation of effective one-body gravitational waveforms Damour:2009zoi, solving the lens equation in strong lensing analyses (see, e.g., Refs. Grespan:2023cpa; Poon:2024zxn), and nuclear reaction networks Reichert:2023xqy.
@@ V Conclusion
Additionally, our framework introduces a novel optimization scheme that recovers the EOS of a given mass-radius or mass-tidal deformability relation by performing gradient descent on a loss function formulated in terms of the EOS parameters, differentiating through the TOV equations. Using this tool, we have discussed a degeneracy present in our metamodel parametrization, which might affect inferences of nuclear empirical parameters from detailed knowledge of NSs using certain EOS parametrizations. This optimization scheme offers a new way to investigate the relationship between the EOS and the properties of NSs, and future work will investigate its robustness when the parametrization of the EOS is misspecified or noisy observations are used.
@@ Appendix A Method validations
In Table 3, we extend the results of Table 2 and show $M_{\rm{TOV}}$, the pressure at $3n_{\rm{sat}}$, $p(3n_{\rm{sat}})$, and the maximal central density $n_{\rm{TOV}}$ at the core of an NS, between Ref. Koehn:2024set and this work for the EOS constraints discussed in Sec. II.5.
Table 3: Comparison of inferred $M_{\rm{TOV}}$, the pressure at $3n_{\rm{sat}}$, $p(3n_{\rm{sat}})$, and the maximal central density $n_{\rm{TOV}}$ at the core of an NS, between Ref. Koehn:2024set and this work for various EOS constraints. Uncertainties are quoted at the $95\%$ credible level.
Constraint $_{\rm{TOV}} [MM_{\odot}$] $p(3n_{\rm{sat}})$ [MeV fm-3] $n_{\rm{TOV}}$ $n_{\rm{sat}}$]
Ref. Koehn:2024set This work Ref. Koehn:2024set This work Ref. Koehn:2024set This work
@@
Radio timing $35_{-0.29}^{+0.73} 2.2$$20_{-0.26}^{+0.39}$ $111_{-49}^{+140}$ $\phantom{0}97_{-49\phantom{0}}^{+65\phantom{0}}$ $5.51_{-1.66}^{+1.89}$ $5.68_{-1.74}^{+1.63}$
@@
All $25_{-0.22}^{+0.42} 2.2$$23_{-0.23}^{+0.41}$ $\phantom{0}90_{-31\phantom{0}}^{+71\phantom{0}}$ $\phantom{0}92_{-42\phantom{0}}^{+70\phantom{0}}$ $5.92_{-1.38}^{+1.35}$ $5.92_{-1.49}^{+1.38}$
Figure 5: Comparison of prior distributions between Ref. Koehn:2024set (Koehn+2024) and this work on the EOS quantities shown in Tab. 2 and Tab. 3. Since both use the EOS parametrization described in Sec. II.1, with priors given by Tab. 1, remaining differences are due to our jax implementation.
@@
Our EOS code and TOV solver are available open source at nuclear-multimessenger-astronomy/jester. The code used in this study is available at Ref. Github_repo_paper.